H!
please help me with this. I dont get this. Thanks for any help coming this way

A can is cylinder for chocolates. The can will contain 75 wrapped chocolates, each with the voulme of 19 cm (cubed). The can has the height of 15 cm. WHat is the radius.

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

A can is cylinder for chocolates. The can will contain 75 wrapped chocolates, each with the volume of 19 cm (cubed). The can has the height of 15 cm. WHat is the radius.
----------
Formula: V = Bh = (pi)r^2*h
19*75 = (pi)r^2*15
(pi)r^2 = 19*15
r^2 = 90.72
r = 9.52 cm
===============
Cheers,
Stan H.

Answer by checkley71(8403)

75*19=1,425 CM^3 IS THE VOLUME OF CANDY.
1,425=14*3.14*R^2
1,425=43.96R^2
R^2=1,425/43.96
R^2=32.42
R=SQRT32.42
R=5.69 CM IS THE RADIUS.